OpenStudy (yanasidlinskiy):

Relations and Functions. Equation: y-6x=4 x: -5, -1, 0, 4 y: Complete the table of values for the equation

OpenStudy (yanasidlinskiy):

My answers are -26, 10, 0, -20 What did I do wrong?

OpenStudy (welshfella):

-26 is right when x = -1 , y -6(-1) = 4 y + 6 = 4 y = 4-6 = -2

OpenStudy (yanasidlinskiy):

Yes, -26 is correct... The last 3 answers are incorrect.

OpenStudy (amoodarya):

y=6x+4 if you put x=-5 ----> y=6(-5)+4=-30+4=-26 if you put x=0 y ?

OpenStudy (welshfella):

yes

rishavraj (rishavraj):

when x = 0 y - 6(0) = 4 y = 4

OpenStudy (welshfella):

y - 6(0) = 4 y - o = 4 y = ?

OpenStudy (yanasidlinskiy):

Ok, yeah, that one I saw after I made the stupid mistake. Ok, I do not see how 10 is wrong!

rishavraj (rishavraj):

when x = -1 y - 6(-1) = 4 y + 6 = 4 y = ??

OpenStudy (yanasidlinskiy):

Oh wait! Guys never mind! I see what I did! These were just complete stupid mistakes at 7:20am!

OpenStudy (yanasidlinskiy):

Thanks for the help!

OpenStudy (welshfella):

y + 6 = 4 subtract 6 from both sides y + 6 - 6 = 4 - 6 y = 4-6